OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y? A) x = 8, y = 4 B) x = 10, y = 2 C) x = 18, y = -6 D) x = 20, y = -8 @kelliegirl33 @maxdrizner

OpenStudy (anonymous):

B

OpenStudy (anonymous):

10-2*2=10-4=6 10+2=12

OpenStudy (anonymous):

Okay, thanks. I'm going to wait until kelliegirl33 is done typing :).

OpenStudy (anonymous):

x + y = 12 --->(-1) x - 2y = 6 --------- -x - y = -12 (result of multiplying by -1) x - 2y = 6 ----------add -3y = -6 y = -6/-3 y = 2 x + y = 12 x + 2 = 12 x = 12 - 2 x = 10 answer B......I agree with @L-Lawliet-L

OpenStudy (anonymous):

Ohhh I see now :D. Thanks @kelliegirl33 for the thorough explanation, once again. Big big big help :). Also, thanks to you @L-Lawliet-L for assisting me too.

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (anonymous):

You are welcome.